OpenStudy (anonymous):

Find the average value of the function over the given interval. (Round your answer to four decimal places.) f(x) = 4 − x2, [−2, 2]

OpenStudy (compassionate):

@Hero , @Satellite7 , @kropot72

OpenStudy (zzr0ck3r):

The average value of \(f\) on the interval \([a,b]\) is given by \[\frac{1}{b-a}\int_a^bf(x)dx\]

OpenStudy (zzr0ck3r):

What is the integral of the function?

OpenStudy (zzr0ck3r):

without the constant...

OpenStudy (anonymous):

yeah I've gotten the same answer like a million times but it says its wrong the integral is (-2,2)

OpenStudy (anonymous):

why did you switch 4x & \[x ^{3}\]

OpenStudy (zzr0ck3r):

\[\dfrac{1}{2-(-2)}[4x-\frac{1}{3}x^3]_{-2}^2\]

OpenStudy (zzr0ck3r):

sorry brain old

OpenStudy (anonymous):

okay so yeah i did that.. will you calculate it for me? see what you get?

OpenStudy (anonymous):

I got 2.6665

OpenStudy (anonymous):

OH MY WORD hahaha there's a part b though too

OpenStudy (anonymous):

Just kidding, I got it :)